OpenStudy (anonymous):

Grizzly the cat (mass = 7.0 kg) was walking along at 1.3 m/s, when out of nowhere a vicious dog called Cedric (poodle) started to chase Griz. If Griz increased her speed to 10.3 m/s in 0.5 s, determine the force exerted by Griz in the acceleration.

OpenStudy (anonymous):

m = 7.0 kg v(initial) = 1.3 m/s v(final) = 10.3 m/s t = 0.5 s \[a = \Delta v / \Delta t\]so a = (10.3-1.3)/0.5 = 9/0.5 = 18 m/s^2 \[F = ma\]so F = (7)(18) = 126 N Answer: F = 126 N

OpenStudy (anonymous):

thx so much